# OpenFACADES: An Open Framework for Architectural Caption and Attribute Data Enrichment via Street View Imagery

### v1

@@ 4.2.1 General performance
As validation, we benchmark our fine-tuned VLMs against various baselines, including ChatGPT-4o and the InternVL2.5 family of models (zero-shot), to evaluate their effectiveness in building-related tasks. The same hyperparameters are applied to fine-tune each VLM, with the number of epochs set to 3 and a learning rate of 8e-6 in a full fine-tuning setup. The results, consolidated in Tables 4(a) and 4(b), confirm the advantages of employing VLMs in these diverse tasks by matching or outperforming ChatGPT-4o baselines in most scenarios.
Table 4: Validation performance comparison of large language models on prediction tasks: (a) building type and surface material; (b) building floors and building age.
@@
Attribute Model Size Accuracy Precision Recall F1 Acc@2
@@
InternVL2.5 1B 0.70 29.22 0.64 52.35
2B 0.70 29.24 0.66 52.03
4B 0.71 29.11 0.64 51.85
In classification tasks, building type and surface material performances are assessed using accuracy, F1-score, recall, and precision. Fine-tuned VLMs exhibit competitive performance, achieving a Top-2 classification accuracy of 75% for building type and 82% for surface material. For predicting the number of building floors and building age, we employ R-squared, Mean Absolute Error (MAE), Mean Absolute Percentage Error (MAPE), and Root Mean Squared Error (RMSE) as evaluation metrics. In building floor prediction, the fine-tuned InternVL2.5 series of models achieve a high R-squared value of approximately 0.77, an MAE of 2.3 floors, and an RMSE of around 4.5 floors, demonstrating substantial improvements over zero-shot baselines. Similarly, in building age prediction, InternVL2.5 models deliver strong results, with an R-squared of approximately 0.70 and an MAE of 29 years, highlighting their effectiveness in capturing architectural and temporal characteristics.
Table 5: METEOR and ROUGE-L evaluation in zero-shot and fine-tuned settings.
Model Size METEOR ROUGE-L
@@
InternVL2.5 1B 39.05 39.06
2B 40.10 39.76
4B 39.17 39.27
Furthermore, the text generated by ChatGPT-4o is used as a baseline to evaluate performance gains in image captioning. In this study, we adopt the METEOR and ROUGE_L metrics, as they assess semantic similarity and linguistic coherence, making them more suitable for evaluating architectural characteristics. Table 5 presents the evaluation metrics across different models. Similar to the labeling task, pre-trained models demonstrate captioning capabilities compared to ChatGPT-4o, while fine-tuned models acquire additional domain-specific knowledge. We observe that increasing model size generally improves performance; however, once the parameter count doubles, marginal gains diminish. This effect is likely due to inherent noise in OSM-derived ground-truth labels and the limited scale of the dataset, which may reduce the advantages of larger architectures. Balancing computational cost with predictive accuracy, we select the 2B-parameter model for subsequent experiments, as it allows efficient inference on a single 24 GB GPU while still delivering notable performance improvements.
4.2.2 Performance by cities
Figure 8 presents the performance of three model variants — (1) the InternVL2.5‐2B model before fine‐tuning and (2) after fine‐tuning, as well as (3) a ChatGPT‐4o reference baseline — on seven cities and four building attributes. In general, the fine‐tuned InternVL2.5‐2B outperforms its non‐fine‐tuned counterpart, showing consistent gains in classification accuracy (Acc.) for building type and surface material, as well as higher R-squared for predicting building floors and age. These improvements are particularly notable in Berlin and San Francisco, where building material, floor and age performance improve substantially. Amsterdam and Helsinki also exhibit moderate but still positive gains for different tasks.
Despite the overall upward trend, improvement magnitude varies across cities and attributes, which may due to several reasons. First, the availability of diverse and distinct samples plays a crucial role: cities with a richer variety of building facades (e.g., Amsterdam, Berlin) yield more pronounced performance boosts. Conversely, locations with more homogeneous or ambiguous building styles (e.g., Manila, Helsinki) show relatively smaller gains. Second, crowdsourced labels in certain cities may be incorrect or insufficient, which can adversely affect the model’s ability to learn reliable city‐specific patterns, restraining potential performance gains.
Nevertheless, when benchmarked against the ChatGPT‐4o baseline, a robust reference point due to its extensive multimodal pretraining, the fine‐tuned InternVL2.5‐2B model demonstrates generally competitive or superior performance. These results confirm that open‐access vision–language models, exemplified by InternVL2.5‐2B, can achieve near‐state‐of‐the‐art performance at no additional licensing cost once adequately fine‐tuned on relevant datasets. This highlights the effectiveness of VLMs in traditional prediction tasks for multiple building attributes across global cities, providing a cost‐effective solution for a wide range of urban remote sensing applications.
Figure 8: Model performance on building attributes across different cities before and after fine-tuning the InternVL2.5-2B model, compared to the baseline performance of ChatGPT-4o. Building type and surface material are evaluated using classification accuracy (Acc.), while number of floors and building age are assessed using R-squared (R2). “NA” indicates cities with insufficient data for model evaluation (ground‐truth instances fewer than 20 in test set).
4.2.3 Performance by categories
Figure 9 presents the confusion matrices illustrating the performance of our VLM on building type and surface material on different categories. Overall, the model demonstrates robust performance for most classes. In terms of well‐predicted labels, visually distinctive building types such as apartments and houses show consistently high accuracies. These categories often have defining features (e.g., apartment blocks characterized by uniform facades and repetitive windows) that the model effectively captures. Similarly, for surface material, high‐frequency and visually salient classes like brick, wood, and glass yield strong performances. Conversely, certain labels are harder to classify, yielding relatively lower accuracies. For building type, hotel or public categories are frequently misidentified as office, suggesting significant overlap in their architectural appearance (e.g., multi-stories, institutional buildings). Likewise, plaster and concrete exhibit misclassifications due to shared grayscale tones and blank textures.
Figure 9: Confusion matrices illustrating the performance of the InternVL2.5-2B model on classifying different categories of building type and surface material. Darker cells indicate higher prediction accuracy.
Figure 10 illustrates the model’s ability to predict building floors and building age under two evaluation schemes: detailed class matching (left) and general range matching (right). For number of floors, the confusion values indicate a good performance for various classes, while when the number increases beyond four or five, misclassification becomes more pronounced. In particular, tall buildings tend to overlap with adjacent categories, highlighting the difficulty of accurately distinguishing high-rise structures based solely on external features and single observation points. When evaluated under the general range scheme, performance improves significantly, suggesting that the model is capable of capturing overall floor patterns from images.
Figure 10: Confusion matrices illustrating the performance of the InternVL2.5-2B model on predicting the number of floors and building age, evaluated based on the accuracy of predictions falling within specific (left) and general (right) ranges. Darker cells indicate higher prediction accuracy.
For building age, the left matrix (exact matching) indicates that categories representing buildings constructed after 1900 exhibit comparatively better performance. This is likely due to their more distinctive architectural styles, which the model can more easily differentiate. In contrast, older buildings (e.g., before 1700, 1800–1900) show greater confusion both among themselves and with intermediate categories (e.g., 1900–1950). This may be attributed to subtle external differences among buildings from these periods and the impact of renovations, additions, or retro designs, which can obscure their original architectural cues (Sun et al., 2022b). Furthermore, the model’s performance may be also influenced by the limited representation of these older eras in the original VLM training dataset.
Taken together, these confusion matrices suggest that the model is capable of inferring the number of floors and the era of construction. However, inherent visual ambiguities — particularly among structurally or stylistically similar categories — contribute to overlaps in predictions. These confusions often arise from overlapping visual cues, which may be further exacerbated by dataset biases, such as incorrect labeling, insufficient sample representation for certain categories, or poor observation. Enhancing the quality, diversity, and coverage of crowdsourced data would be a valuable step toward improving the dataset and the model’s performance.
4.2.4 Ablation experiments on data size
@@
Figure 11 presents the results of ablation experiments on the InternVL2.5-2B using different combinations of a training set from 58,942 image-text pairs. The left panel shows accuracy for material and type predictions, while the right panel illustrates R-squared for floor and age predictions across varying dataset percentages. Table 6 summarizes semantic similarity and linguistic coherence relative to GPT‐generated captions under different conditions.
@@
These ablation experiments reveal several important insights. First, in multi-attribute prediction tasks (Figure 11), performance peaks early in both scenarios of adjusting either OSM data or GPT-generated data. Even smaller datasets (around 5–10% of the full corpus) yield notable performance gains, highlighting the VLM’s ability to learn effectively in data-constrained scenarios. This behavior can be attributed to the pre-trained semantic relationships embedded in the VLM’s latent space from its foundational training. Fine-tuning on limited data stabilizes outputs by aligning task-specific features with the model’s pre-existing knowledge distribution. Second, performance rises gradually when adding OSM ground truth data for most attribute prediction tasks, while GPT-generated data slightly diminishes performance gains. One plausible explanation is that OSM data encodes structured, human‐validated geographic knowledge, whereas GPT‐generated samples may introduce inaccuracies or hallucinated features that misalign the VLM’s learned representations. Mitigating such noise — by refining annotation procedures or excluding low‐quality samples — could improve overall accuracy and robustness (Chen et al., 2024b). Third, adding OSM data constrains the descriptive ability across tasks (Table 6). This trade-off may reflect task interference in multi-task learning: optimizing for structured building attributes could suppress the model’s capacity to generate diverse captions. Addressing this may involve upscaling model capacity, curating high‐quality OSM–GPT hybrid datasets, or leveraging techniques such as knowledge distillation to balance structure with generative expressiveness.
Table 6: METEOR and ROUGE-L evaluation across OSM and GPT splits.
Dataset METEOR ROUGE-L

### v2

@@ 5.2.1 General performance
After we fine-tuned the baseline models mentioned in Section 4.2, we compare their performance with that of the zero-shot ChatGPT and fine‐tuned InternVL3‐2B on four building characteristics, as summarized in Tables 5(a) and 5(b). To facilitate a fair comparison, we integrate GPT-generated data to supplement the missing OSM data in the training set for the CV models.
Table 5: Validation performance comparison among zero-shot ChatGPT, fine-tuned InternVL3-2B and CV models.
Attribute Model LR Epoch Acc (%) mPre mRec mF1 Acc@2 (%)
Building type ChatGPT (zero-shot) - - 57.69 0.649 0.566 0.571 75.06
DenseNet 1e-5 14 53.42 0.540 0.529 0.529 69.39
VGG 5e-4 9 47.66 0.492 0.470 0.462 62.03
ResNet18 1e-4 19 50.59 0.509 0.499 0.493 66.28
ResNet50 1e-4 24 52.20 0.526 0.512 0.504 65.98
ResNet101 5e-5 13 53.58 0.543 0.527 0.524 68.52
ViT16 5e-6 7 53.60 0.549 0.532 0.530 69.35
ViT32 1e-5 5 52.24 0.519 0.515 0.507 67.93
Swin_b 5e-6 17 56.11 0.562 0.558 0.554 73.14
InternVL3-2B 8e-6 3 61.27 0.661 0.602 0.609 77.31
Surface material ChatGPT (zero-shot) - - 65.41 0.581 0.614 0.553 79.07
DenseNet 1e-4 5 65.92 0.575 0.610 0.566 81.45
VGG 1e-5 12 57.02 0.504 0.529 0.494 74.06
ResNet18 1e-5 23 63.78 0.560 0.587 0.554 79.57
ResNet50 5e-5 30 65.79 0.601 0.610 0.578 81.08
ResNet101 5e-5 25 66.67 0.608 0.627 0.582 79.70
ViT16 5e-5 18 64.79 0.571 0.587 0.564 81.20
ViT32 1e-5 13 60.15 0.538 0.554 0.511 77.07
Swin_b 5e-6 33 69.17 0.623 0.655 0.612 85.84
InternVL3-2B 8e-6 3 68.05 0.598 0.634 0.588 81.70
Attribute Model LR Epoch R2 ($\uparrow$) MAE ($\downarrow$) MAPE (%) ($\downarrow$) RMSE ($\downarrow$)
Number of floors ChatGPT (zero-shot) - - 0.721 2.36 38.66 5.01
DenseNet 5e-5 5 0.774 2.53 41.91 4.51
VGG 5e-5 12 0.689 3.10 42.43 5.28
ResNet18 5e-5 16 0.741 2.82 38.81 4.82
ResNet50 5e-5 23 0.768 2.48 41.74 4.56
ResNet101 5e-5 14 0.777 2.48 41.97 4.48
ViT16 5e-6 6 0.765 2.63 46.47 4.59
ViT32 1e-6 22 0.727 2.83 45.08 4.95
Swin_b 1e-5 10 0.773 2.41 41.48 4.52
InternVL3-2B 8e-6 3 0.789 2.13 36.93 4.35
Building age ChatGPT (zero-shot) - - 0.645 31.63 58.79 57.07
DenseNet 1e-5 21 0.720 31.31 78.36 50.68
VGG 1e-4 7 0.559 42.55 121.31 63.59
ResNet18 5e-5 23 0.714 32.63 74.44 51.23
ResNet50 1e-4 18 0.721 32.26 85.92 50.55
ResNet101 5e-5 18 0.738 30.61 78.77 49.01
ViT16 5e-6 29 0.719 30.89 80.02 50.77
ViT32 1e-5 23 0.675 33.82 71.89 54.61
Swin_b 5e-6 27 0.723 32.72 93.77 50.40
InternVL3-2B 8e-6 3 0.710 28.06 58.27 51.50
The fine-tuned InternVL3-2B achieves the highest overall performance, particularly in predicting building type and number of floors. Although CV models slightly outperform the VLM in surface material classification and building age estimation, they require attribute-specific architectures and domain-specific tuning. In contrast, VLMs provide a unified and adaptable framework, delivering comparable or superior results across multiple prediction tasks. Notably, zero-shot ChatGPT also demonstrates strong capability across all tasks, achieving performance close to that of some fine-tuned CV models. This suggests its potential as a practical tool for supplementing building data when labeled samples are limited. This advantage of VLM can be attributed primarily to the semantic reasoning and contextual understanding inherent in pretrained VLMs, which significantly enhance their generalization across diverse tasks. Moreover, VLMs can simultaneously infer multiple target variables and implicitly model correlations among tokens associated with different tasks, whereas the CV baselines considered here are independently fine-tuned for each attribute. The combination of robust predictive performance and enhanced functionality makes VLMs a compelling alternative to traditional approaches in future applications.
5.2.2 Performance by cities
Figure 9 presents the performance of three model variants: (1) the InternVL3‐2B model before fine‐tuning and (2) after fine‐tuning, as well as (3) a ChatGPT‐4o reference baseline, on seven cities and four building attributes. In general, the fine‐tuned InternVL3‐2B outperforms its non‐fine‐tuned counterpart, showing consistent gains in predicting all four attributes. These improvements are particularly notable in Berlin and San Francisco, where building type, material, and floor performance improve substantially. Amsterdam and Helsinki also exhibit moderate but still positive gains for different tasks.
Despite the overall upward trend, improvement magnitude varies across cities and attributes, which may due to several reasons. First, the availability of diverse and distinct samples plays a crucial role: cities with a richer variety of building facades (e.g., Amsterdam, Berlin) yield more pronounced performance boosts. Conversely, location with more homogeneous or ambiguous building styles (e.g., Manila) shows relatively smaller gains. Second, crowdsourced labels in certain cities may be incorrect or insufficient, which can adversely affect the model’s ability to learn reliable city‐specific patterns, restraining potential performance gains.
Nevertheless, when benchmarked against the zero-shot ChatGPT‐4o baseline, the fine‐tuned InternVL3‐2B model demonstrates generally competitive or superior performance. These results confirm that open‐access VLMs can achieve near‐state‐of‐the‐art performance at no additional licensing cost once adequately fine‐tuned on relevant datasets. This highlights the effectiveness of VLMs in predicting multiple building attributes across global cities, providing a cost‐effective solution for a wide range of urban remote sensing applications.
Figure 9: Model performance on building attributes across different cities before and after fine-tuning the InternVL3-2B model, compared to the baseline performance of ChatGPT-4o. Building type and surface material are evaluated using classification accuracy (Acc.), while number of floors and building age are assessed using R-squared (R2). “NA” indicates cities with insufficient data for model evaluation (ground‐truth instances fewer than 20 in test set).
5.2.3 Performance by categories
Figure 10 presents the confusion matrices illustrating the performance of our VLM on building type and surface material on different categorical labels. Overall, the model demonstrates robust performance for most classes. In terms of well‐predicted labels, visually distinctive building types such as apartments and houses show consistently high accuracies. These categories often have defining features (e.g., apartment blocks characterized by uniform facades and repetitive windows) that the model effectively captures. Similarly, for surface material, high‐frequency and visually salient classes like brick, wood, and glass yield strong performances. Conversely, certain labels are harder to classify, yielding relatively lower accuracies. For building type, hotel or public categories are frequently misidentified as office, suggesting significant overlap in their institutional architectural appearance (e.g., multi-stories, ordered elements). Likewise, plaster and concrete exhibit misclassifications due to shared grayscale tones and blank textures.
Figure 10: Confusion matrices illustrating the performance of the InternVL3-2B model on classifying different categories of building type and surface material. Darker cells indicate higher prediction accuracy (%).
Figure 11 illustrates the model’s ability to predict the number of floors and building age. The heatmaps provide a detailed visualization of prediction accuracy relative to ground-truth values, with darker colors indicating higher frequencies of accurate predictions within each bin. For the number of floors (left), the heatmap demonstrates strong prediction performance for lower-rise buildings. As the number of floors increases beyond five, prediction performance start to decline. High-rise structures (stories beyond 15) exhibit more frequent errors, highlighting the challenge in distinguishing floor count for taller buildings from single viewpoint images.
For building age (right), the heatmap reveals stronger predictive performance for more recent buildings (post-1900), reflecting their typically more distinctive and recognizable architectural characteristics. Conversely, older buildings (pre-1900) exhibit larger errors, with significant overlaps between adjacent historical periods. The subtle external distinctions, combined with modifications such as renovations and retrofits, contribute to the difficulty in accurately classifying older structures based solely on visual appearance (Sun et al., 2022b). Additionally, the model’s performance is likely influenced by the limited representation of older building periods within the original training dataset. Nevertheless, the overall R-squared above 0.7 confirms the model’s capability in capturing broad temporal patterns of floor count and building age from images.
Figure 11: Heatmaps illustrating the performance of the InternVL3-2B model in predicting the number of floors (left) and building age (right). Darker colors indicate a higher frequency of predictions within each ground-truth bin.
Taken together, these results suggest that the model is capable of inferring attributes across various ranges. However, inherent visual ambiguities, particularly among structurally or stylistically similar categories, contribute to overlaps in predictions. Enhancing the quality, diversity, and coverage of crowdsourced data would be a valuable step toward improving the dataset and the model’s performance.
5.3 Generalizability and robustness
Generalization to unseen city.
CV models with good performance are selected for experiments in this section. Tables 6(a) and 6(b) show that the VLM model exhibits superior generalizability compared to commonly used computer vision models, particularly in the tasks of building type and age prediction. This improved performance can be attributed to the VLM’s ability to leverage semantic reasoning and contextual understanding from large-scale image-text pretraining, allowing it to adapt more effectively to diverse architectural features. While the performance on number of floors prediction is comparable across models, the lower accuracy of Swin_b on other attributes suggests a reliance on localized visual features, which limits its generalization capability. Despite these strengths, the VLM’s low macro-F1 score for building-type classification highlights persistent challenges, notably the underrepresentation of minority classes and ambiguity in building features. To address these issues, assembling a more balanced and diverse building imagery dataset and incorporating domain-specific fine-tuning are recommended. Collectively, these findings underscore the potential of VLMs to enhance robustness and generalizability across varied urban contexts.
(a) Performance on classification task of building type
Attribute Model Acc (%) mPre mRec mF1
Building type DenseNet 31.12 0.25 0.33 0.21
ResNet101 31.20 0.24 0.34 0.21
ViT16 32.54 0.26 0.32 0.21
Swin_b 24.58 0.26 0.33 0.20
InternVL3-2B 63.80 0.44 0.50 0.42
Attribute Model RMSE ($\downarrow$) MAE ($\downarrow$) MAPE (%) ($\downarrow$) R2 ($\uparrow$)
Number of floors DenseNet 1.82 1.13 30.20 0.610
ResNet101 1.90 1.15 33.46 0.576
ViT16 1.79 1.15 33.10 0.624
Swin_b 1.74 1.08 29.39 0.645
InternVL3-2B 1.73 0.90 26.20 0.645
Building age DenseNet 55.50 43.22 74.89 0.200
ResNet101 51.55 39.83 172.65 0.166
ViT16 51.44 38.85 168.38 0.157
Swin_b 54.92 42.27 195.16 0.170
InternVL3-2B 44.67 30.76 66.66 0.600
Robustness to varying image quality.
As described in Section 4.3, we evaluate the robustness of the VLM against image corruptions by testing it on a perturbed dataset derived from the test set. Figure 12 illustrates the model’s performance under varying severity levels of occlusion, motion blur, Gaussian noise, and brightness distortions. In general, the model demonstrates resilience, with performance dropping by less than 10% under most mild and moderate image corruptions. In particular, the model remains significantly stable in handling lighting variations and occlusion, both of which are common challenges in crowdsourced image datasets. However, the model experiences a significant performance drop when confronted with moderate to severe noise and blurriness. In particular, the model is most affected by motion blur, where the error rate for the number of floors prediction increases from 0.21 (clean error) to 0.48, and the error rate for building age prediction rises from 0.29 (clean error) to 1.45. These findings emphasize the necessity of image preprocessing techniques to filter out degraded images during the image selection stage.
Furthermore, the Relative mmCE is computed for different models using ResNet50 as the baseline. Table 7 presents the relative error rates across different building attribute prediction tasks, indicating each model’s stability compared to the baseline. In general, different models demonstrate various capability in handing corruption data. Multi-attribute prediction VLM (InternVL3-2B) demonstrates superior stability compared to single-attribute CV models in most cases, having lowest mean Relative mmCE among all attributes. Especially, when distinguishing building type and building age when encountering occlusion, motion and brightness variations, VLM demostrate well stability. CNN models’ stability performs comparably to more advanced models in the tasks of number of floors prediction, while ViT model performs superior in handling data with Gaussian noise. This outcome implies that additional domain-specific constraints or specialized training strategies might be required to enhance performance on crowdsourced image data.
Table 7: Relative prediction error compared to ResNet50 under different corruptions and perturbations for different objective attributes.
Attribute Model Occlusion Motion Noise Brightness Relative mCE
Building type ResNet50 1.00 1.00 1.00 1.00 1.00
DenseNet 2.10 0.98 0.86 0.78 1.18
ResNet101 0.89 0.95 0.61 0.54 0.75
ViT16 0.76 0.83 0.29 1.22 0.78
Swin_b 0.88 1.12 0.30 0.47 0.69
InternVL3-2B 0.51 0.42 0.39 0.46 0.44
Surface material ResNet50 1.00 1.00 1.00 1.00 1.00
DenseNet 1.26 1.09 1.20 1.16 1.18
ResNet101 1.49 0.98 1.25 1.02 1.19
ViT16 0.48 0.72 0.50 1.48 0.80
Swin_b 0.99 1.11 1.03 0.70 0.96
InternVL3-2B 0.59 0.66 0.64 0.47 0.59
Number of floors ResNet50 1.00 1.00 1.00 1.00 1.00
DenseNet 2.01 1.62 0.86 1.24 1.43
ResNet101 0.80 1.19 0.31 0.69 0.75
ViT16 1.07 1.41 0.46 1.61 1.14
Swin_b 0.66 1.87 0.49 0.84 0.97
InternVL3-2B 0.63 0.78 0.42 0.54 0.59
Building age ResNet50 1.00 1.00 1.00 1.00 1.00
DenseNet 2.65 0.94 0.99 1.43 1.50
ResNet101 1.04 0.78 0.48 1.10 0.85
ViT16 0.50 0.78 0.27 1.80 0.84
Swin_b 0.67 1.11 2.90 1.43 1.53
InternVL3-2B 0.30 0.69 0.48 0.54 0.50
In conclusion, building on insights from CV baselines, VLMs not only demonstrate robust and generalizable features for tackling diverse tasks based on crowdsourced data, but they also represent a promising framework for large-scale or cross-regional implementations that demand multi-feature prediction and flexible adaptation.
5.4 Ablation experiments
Fine-tuning settings.
As shown in Table 8, we compare model performance across different settings, including LoRA, full fine-tuning, number of epochs, and learning rate. We summarize overall performance using two aggregate metrics: Overall Accuracy (OA), computed as the average of accuracies, and Overall R2 (OR), computed as the average R2. In general, full fine-tuning consistently outperforms LoRA across most tasks. Although LoRA offers parameter efficiency and retains the original checkpoint’s general capabilities, it typically yields lower performance. Therefore, we adopt full fine-tuning for further investigation.
We further investigated the effect of training epochs under the default learning rate of 4e-5. Performance generally improves from epoch 1 to 3, with diminishing or unstable gains beyond that point. For example, at epoch 3, full fine-tuning achieves strong and balanced results, including a peak OA of 64.02% and competitive OR. Although epoch 4 yields a slightly higher OR (0.725), the improvements are marginal and come at the cost of additional computation. Thus, epoch 3 is selected as the most efficient convergence point.
We also experimented with several learning rates, including the default 4e-5 and smaller values such as 4e-6. A learning rate of 8e-6 offers improved stability and slightly better OR. In addition, our observations suggest that smaller learning rates are especially beneficial for fine-tuning on relatively small datasets, as they help retain general-purpose capabilities such as open-ended reasoning and semantic alignment. Based on these comparisons, we adopt full fine-tuning with 3 training epochs and a learning rate of 8e-6 as our final configuration. This setting provides the best trade-off between predictive performance, training efficiency, and the preservation of the pretrained strengths of large VLMs.
Table 8: Validation performance across epochs and learning rates for LoRA and full‐tuning schemes.
Epoch LR Acc (%) mF1 OA R2 MAE OR
Type Material Type Material Floor Age Floor Age
LoRA
1 4e-5 60.29 60.29 0.583 0.513 60.29 0.739 0.524 2.64 40.44 0.632
2 4e-5 60.66 62.50 0.593 0.532 61.58 0.729 0.560 2.60 38.01 0.645
3 4e-5 60.52 63.24 0.592 0.537 61.88 0.745 0.595 2.52 35.79 0.670
4 4e-5 60.11 64.34 0.583 0.548 62.22 0.741 0.593 2.52 35.73 0.667
5 4e-5 60.34 64.86 0.586 0.553 62.60 0.757 0.571 2.51 35.97 0.664
Full‐tuning
1 4e-5 62.64 63.97 0.608 0.538 63.31 0.719 0.637 2.50 33.25 0.678
2 4e-5 62.02 64.71 0.599 0.542 63.37 0.735 0.672 2.40 30.95 0.704
3 4e-5 63.32 64.71 0.618 0.542 64.02 0.734 0.680 2.38 30.44 0.707
4 4e-5 62.09 65.07 0.605 0.545 63.58 0.748 0.701 2.31 29.03 0.725
5 4e-5 62.41 64.34 0.608 0.541 63.38 0.741 0.696 2.39 29.14 0.719
3 4e-4 10.31 16.54 0.030 0.085 13.43 -0.291 -0.060 7.88 79.28 -0.176
3 8e-5 62.36 63.24 0.609 0.546 62.80 0.736 0.646 2.43 31.73 0.691
3 4e-5 63.32 64.71 0.618 0.542 64.02 0.734 0.680 2.38 30.44 0.707
3 8e-6 63.43 63.97 0.619 0.540 63.70 0.762 0.665 2.19 31.63 0.714
3 4e-6 63.39 62.50 0.620 0.532 62.94 0.748 0.653 2.33 31.91 0.701
Model size.
We further compare the performance of VLM variants with different model sizes under both zero-shot and fine-tuned settings, as shown in Tables 9(a) and 9(b). In the zero-shot setting, larger model sizes or more recent pretrained weights generally yield better predictions.
The results confirm that fine-tuning is essential for domain adaptation, consistently improving performance across all building profiling tasks. For example, classification accuracy improves by 7–15 percentage points over zero-shot baselines, while R2 values for floor and age prediction also increase significantly. After fine-tuning, model performance becomes more comparable across sizes, though task-specific variations remain. InternVL2.5–4B achieves the highest accuracy in classifying building type and surface material, whereas InternVL3 variants perform better on floor count and age prediction. These findings suggest that models in the 1–2B parameter range are generally sufficient for building profiling, particularly when training on relatively small datasets.
Table 9: Validation performance of VLM variants with different model sizes under both zero-shot and fine-tuning conditions.
@@ 5.4 Ablation experiments
Attribute Model Size Acc (%) mPre mRec mF1 Acc@2 (%)
@@
InternVL3 1B 0.713 28.09 58.15 51.36
2B 0.710 28.05 58.27 51.50
InternVL2.5 4B 0.707 29.11 63.73 51.85
Table 10 presents captioning evaluation metrics for various models. As with the labeling task, pretrained models already exhibit reasonable performance relative to the ChatGPT-4o reference captions, while fine-tuned models incorporate domain-specific knowledge and produce more coherent, better-structured captions. We also observe that performance generally improves with model size, but doubling the parameter count yields diminishing marginal gains, which is likely a consequence of the dataset’s limited scale and the inherent noise in OSM derived ground truth labels. To balance computational cost and accuracy, we therefore select the InternVL3 model with 2 billion parameters for this study.
Table 10: Captioning performance metrics (%) in zero-shot and fine-tuned settings, evaluated against ChatGPT-4o-generated reference captions.
Model Size BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L
@@
InternVL3 1B 56.60 41.62 32.65 25.99 44.46 43.31
InternVL3 2B 56.91 41.91 32.93 26.25 44.91 43.55
InternVL2.5 4B 52.32 37.68 29.00 22.64 41.02 40.43
Data size.
@@
To assess the impact of training set size on performance, we trained InternVL3-2B on varying fractions of the 58,942 image–text pairs listed in Table 4. The left panel of Figure 13 reports accuracy for material and type predictions, while the right panel shows R2 values for floor and age predictions at each dataset proportion. Table 11 presents BLEU, METEOR, and ROUGE-L scores relative to GPT-generated captions under each condition.
@@
These ablation experiments reveal several important insights. First, in multi-attribute prediction tasks (Figure 13), performance peaks early in both scenarios of adjusting either OSM data (single-attribute Q&A) or GPT-generated data (multi-attribute Q&A and image captions). Even smaller datasets (around 5–10% of the full corpus) yield notable performance gains, highlighting the VLM’s ability to learn effectively in data-constrained scenarios. This behavior can be attributed to the pre-trained semantic relationships embedded in the VLM’s latent space from its foundational training. Fine-tuning on limited data stabilizes outputs by aligning task-specific features with the model’s pre-existing knowledge distribution. Second, performance rises gradually when adding OSM ground truth data for most attribute prediction tasks, while GPT-generated data slightly diminishes performance gains. One possible explanation is that OSM data encodes structured, human‐validated geographic knowledge, whereas GPT‐generated samples may introduce inaccuracies or hallucinated features. Mitigating such noise by refining annotation procedures or excluding low‐quality samples could improve overall accuracy and robustness (Chen et al., 2024b). Third, integrating single-attribute Q&A data derived from OSM labels appears to constrain descriptive richness across tasks (Table 11). This trade-off likely reflects task interference in a multi-task learning setup, where optimizing for structured attribute prediction can suppress the model’s ability to generate diverse captions. To address this limitation, one could increase model capacity, curate high-quality OSM–GPT hybrid datasets, or employ techniques such as knowledge distillation to balance structured output with generative expressiveness.
Table 11: BLEU, METEOR, and ROUGE-L evaluation across OSM and GPT splits.
Dataset BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L